Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

Wilkinstown Target Shooting Club Open

Distribution

Dr. Trevor West DUCAC Chairman
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Tony O’Donnell Student Record

Summary

Venue Wilkinstown Target Shooting Club
Date 14 October 2001
Discipline 10m Air Rifle

DURC Attendees

Steven Watterson Mark Dennehy
Ben Jones  Peter Byrne

Scores

Steven Watterson 571
Mark Dennehy 544
Ben Jones 485
Peter Byrne 453

Prizes Won

Steven Watterson 3rd Place, Class A

3rd Place Overall Personal Best Score

Mark Dennehy 2nd Place, Class B

Personal Best Score

Peter Byrne 1st Place, Class C

Report

DURC sent a team of four to this year’s WTSC Open, the first competition of the Air Rifle season. Attendance was low due to difficulties with transport and lack of training time for new club members. The UCD Open in two weeks should see a much higher attendance from DURC.

The new WTSC mobile range was used: this range is currently the largest in the country with 15 firing points. It currently hosts the National Squad training sessions as well as other large competitions. Previously, only the UCD range could cope with these events.

The DURC contingent arrived in WTSC early on Sunday morning, despite having attended the Olympic Council of Ireland’s seminar on coaching in the University of Limerick the day before. Mark, Peter and Ben shot on the first detail, Steven acting as “buddy” for all three. The “buddy” system, recently introduced in DURC by Steven, has proven a great success - the “buddy” acts as another pair of hands for the shooter while setting up on his/her firing point and general dogsbody during the match to minimize disruption for the shooter, who then trades role with the “buddy” on the next detail. This allows for one-to-one supervision of new shooters as well as improved performance of experienced shooters.

Steven and Mark both set personal bests of 571 and 544 respectively and both went through to the ten-shot final, where Steven had the best final score of the match. Mark was shooting with his new air rifle and other gear - and despite some silly technical errors that lost him up to 15 points, shot a PB. Proving that the right equipment can make a serious difference to performance. For reference, 565 is the Minimum Qualifying Score, or MQS, for the Olympics; 537 is the qualifying score for the National Developmental Squad; and 559 is the qualifying score for the National Squad.

Peter Byrne and Ben Jones had dissappointing scores given their recent averages - Ben has competed in the finals of the Centenary Bisley postal in air rifle (an international match where he placed fourth out of hundreds), and Peter is currently one of the top three junior shooters in Ireland.

Steven Watterson, DURC Captain at the WTSC Open

Misé le meas,

Mark Dennehy, DURC Secretary